10 bulbs of 50 Watt each glows 10 hours a day. then how much electrical energy is consumed in 30 days:1) 5 kilowatt hour2) 10 kilowatt-hour 3) 150 kilowatt hour4) 200 kilowatt hour

Grade:12th pass

1 Answers

Priya
6 years ago
Energy= Power × timeEnergy consumed per day by 1 bulb=50×10=500 Watt hourThen by 10 bulb energy consumed=500×10=5KWhSo,the electrical energy consumed in 30 days=5KWh×30= 150 Kilowatt hour.So, correct option is (3).